OpenStudy (anonymous):

can someone help me with this equation please. 1/3y+-1/2<= 5/6y+1/2

OpenStudy (anonymous):

multiply every thing by 1/2. then get one y term on one side and ur done.

OpenStudy (texaschic101):

multiply the entire equation by 6 to get rid of the fractions... (6)1/3y - (6)1/2 <= (6)5/6y + (6)1/2 2y - 3 <= 5y + 3 2y - 5y <= 3 + 3 -3y <= 6 (divide by -3) y >= -2 (when dividing/multiplying by a negative, the sign changes) or... 1/3y - 1/2 <= 5/6y + 1/2 1/3y - 5/6y <= 1/2 + 1/2 2/6y - 5/6y <= 1 - 3/6y <= 1 (divide by -3/6) y >= 1 * -6/3 (change the sign) y > = 2 either way you will get the same answer. It just depends on if you want o deal with fractions or not :)

OpenStudy (anonymous):

Ok, THANK YOU =)

OpenStudy (texaschic101):

anytime :)